A typical solid fertiliser contains the elements N, P and K in the ratio of 15g;30g;15g per 100g of fertiliser. The recommended usage of fertiliser is 14g of fertiliser per dm3 of water.

What is the concentration of nitrogen atoms in this solution?
1
Expert's answer
2017-02-09T06:45:40-0500
Solution:
At first we have to know what concentration is meant in the task. As there are no any special notes, we assume that it is weight %, which shows mass of solute per 100 g of solution and is calculated as (mass of solute, g / mass of solution, g) * 100%.
In our case the total mass of solution is (14 g of fertilizer) + (mass of 1 dm3 of water). The mass of 1 dm3 of water is 1000g.
So the mass of solution is 14 g + 1000 g = 1014 g.

Now we find the mass of nitrogen in this solution. We know that 100g of fertilizer contains 15g of nitrogen. Then 14g of such fertilizer contains (14g * 15 g) / 100 g = 2.1 g of nitrogen.

So our 1014 g of solution contain 2.1 g of nitrogen.
Then the concentration is (2.1 g / 1014 g) * 100% = 0.2%.


Answer:
The concentration of nitrogen atoms in the solution is 0.2%